7. Assume that advertising shifts the demand curve for Coca-Cola to the right along the supply curve which pushes the Coca-Cola price up by 45%. If the old equilibrium price of Coke was $1.33/liter bottle and the old equilibrium quantity is 13,360.0 million liter bottles, the elasticity of Coca-Cola supply is 0.50 and the elasticity of demand is -1.83, what is the new equilibrium quantity demanded of Coca-Cola? What is the new equilibrium quantity supplied?
Answer #1

Ans

Here elasticity of demand =1.83.in economics negative Baile of elasticity is not taken into account

This means as price increases by 1 percent demand decreases by 1.83 percent

Demand falls by 1.83(45) =82.35%

So demand =13360-(82.35%)of 13360=2358 million

Similarly supply=13360+0.5(45)% of 13360=16366

Note supply rises as price rises